40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 16 02:27:56 2013, Jenne, Update, LSC, PRMI + 2 arm attempt PRCL_OLTF_PRMIir_XARMgreen_16Oct2013.pdfMICH_OLTF_PRMIir_XARMgreen_16Oct2013.pdfTimeSeries_PRMI_2arms_16Oct2013.png
    Reply  Wed Oct 16 02:39:24 2013, rana, Update, LSC, PRMI + 2 arm attempt 
    Reply  Wed Oct 16 17:34:28 2013, Jenne, Update, LSC, PRMI + 2 arm attempt TimeSeries_PRMI_Xarm_16Oct2013.png
Message ID: 9243     Entry time: Wed Oct 16 02:27:56 2013     Reply to this: 9244   9247
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PRMI + 2 arm attempt 
[Masayuki, Jenne]
 
Masayuki informed me that the Xarm ALS was feeling pretty good today, so we quickly (<20 minutes, including 2 open loop transfer functions) locked the PRMI+Xarm
We then tried PRMI + 2 arms, but while trying to bring the arms into IR resonance, the PRMI lost lock.
 
What we did (procedure-wise):
 
I locked and aligned both arms in IR.  I misaligned the ETMs, locked MICH to tweak BS pointing.  I locked PRMI with REFL 165 I&Q, and used the ASS to tweak up the PRM pointing.  I then moved PRM -0.5 units in pitch (after turning off the LSC). 
Masayuki then restored ETMX, locked the Xarm ALS, used his nice new script to find the IR resonance, then we stepped ~5 offset counts away from the resonance.  This is just barely off the IR resonance, since we don't want to cross any sideband cavity resonances.  I then brought the PRM back, and turned on the LSC.  PRMI immediately acquired lock.  Then Masayuki, with a 30 second ramp time, moved us back the ~5 counts until we had Xarm IR resonance! 
After that,
we took 2 open loop transfer functions, one of PRCL and one of MICH.

It was smooth like butter.  Seriously, we decided to give it a try around 12:50am, and by 1:08am we had saved both OLTF .xml files.  During this lock, Xarm ALS beatnote was -14.5 dBm, at 68.9 MHz.
 
After that success, we decided to be bold, and see if we could do PRMI + 2 arms.  I turned off the PRMI's LSC, and misaligned the PRM by -0.5 slider units.  We restored ETMY, and Masayuki got both arms locked with ALS, found the resonances, and then stepped ~5 offset counts away from each resonance.  I restored the PRM, and enabled the LSC.  Once again, the PRMI acquired lock immediately.  However, when I tried to turn on the ASS, I lost lock of the PRMI (but not the arms' ALS).  PRMI did seem noticeably noisier this time than either without any arms, or even with just Xarm.  The POP spot on the camera was wiggling about the same amount that it normally does when the ETMs are misaligned, PRMI is locked, and the ASS is on.  However, the ASS was off, and we were still seeing this angular motion.  Anyhow, I relocked the PRMI, and left the ASS off.  Then, we tried bringing the arms back into resonance for IR.  We should probably automate this process, or do it with a CARM-type loop, so that both come in at the same time.  As it was, Masayuki typed the offset for resonance into one arm (with a 30 second ramp time), then quickly typed in the number for the other arm (again with a 30 second ramp time).  So, the arms weren't exactly in common.  We lost PRMI lock during the scan toward resonance. 
Here is some time series data.  While it's tricky to see much in this plot, here's the command to get it:  ./getdata -s 1065947469 -d 40 -c C1:LSC-POP22_I_ERR_DQ C1:LSC-TRX_OUT_DQ C1:LSC-TRY_OUT_DQ C1:ALS-OFFSETTER1_OUT_DQ C1:ALS-OFFSETTER2_OUT_DQ C1:LSC-MICH_IN1_DQ C1:LSC-PRCL_IN1_DQ   Also, the text files for these traces are in my directory: /users/jenne/PRCL/PRMI_Xarm_ALS_16Oct2013/  During this trial (at least before the lockloss), Xarm was still on the same lock stretch with ALS as before, so -14.5 dBm at 68.9 MHz, and the Yarm was -21.7 dBm at 82.9 MHz.
 
To-do:
To take a step back, we should try bringing in only one arm to IR resonance, then the other, to see if we can isolate what the cause of the lockloss was. 
I need to finish setting up the new sensing matrix measurement stuff, so we can take sensing matrix data throughout this process.
 
 

Plots:

PRCL Open Loop Transfer Function.  PRMI locked on REFL 165 I&Q, Xarm held on IR resonance using ALS, ETMY misaligned:

PRCL_OLTF_PRMIir_XARMgreen_16Oct2013.pdf

MICH Open Loop Transfer Function.  PRMI locked on REFL 165 I&Q, Xarm held on IR resonance using ALS, ETMY misaligned:

MICH_OLTF_PRMIir_XARMgreen_16Oct2013.pdf

 Time series data during our PRMI + 2 arm attempt:

TimeSeries_PRMI_2arms_16Oct2013.png

ELOG V3.1.3-